OpenStudy (anonymous):

Solve the exponential equation. If the answer is not exact, give the answer to four decimal places. See Example 4. (Objective 1. Enter your answers as a comma-separated list.) 5^x^2=6^x

OpenStudy (anonymous):

\[5^{x ^{2}} = 6^{x}\]

OpenStudy (unklerhaukus):

Taking the log to base 6 on both sides of the equation \[\begin{align} \log_6(5^{x^2})&=\log_6(6^x)\\ x^2\log_6(5)&=x\\ x^2\log_6(5)-x&=0\\ x\Big(x\log_6(5)-1\Big)&=0 \end{align}\] A product of two terms equal zero. The two solution can be found by setting one of terms in the product to zero. \[x=0\] or \[x\log_65-1=0\]

OpenStudy (unklerhaukus):

the first solution was trivial, but for the second, you will have to rearrange the equation a little ,and used the change of base formula

OpenStudy (anonymous):

i dont understand why we plugged in base 6

OpenStudy (unklerhaukus):

\[x\log_6(5)-1=0\\ x\log_6(5)=1\\ x=\frac1{\log_6(5)}=\frac1{\log(5)/\log(6)}=\frac{\log(6)}{\log(5)}\] They other way works too (maybe a little simpler) Taking the log to base 5 on both sides of the equation \[ \begin{align} \log_5(5^{x^2})&=\log_5(6^x)\\ x^2&=x\log_5(6)\\ x^2-x\log_5(6)&=0\\ x(x-\log_5(6))&=0\\ \end{align}\] \[x=0\]or\[x-\log_5(6)=0\\x=\log_5(6)\\x=\frac{\log(6)}{\log(5)}\]

OpenStudy (unklerhaukus):

The reason for taking the log to either base 5 or base 6, is to get rid of those exponentials

OpenStudy (unklerhaukus):

logarithm and exponential are kinda like inverses